OpenStudy (anonymous):

multiply (4+√10)(4-√10)

OpenStudy (anonymous):

This is a difference of 2 squares: \[(a+b)(a-b) = a^{2}-b^{2}, a = 4, b = \sqrt{10}\] so just following the formula we get: \[a^{2} - b^{2} = (4)^{2}-(\sqrt{10})^{2}= 16- 10 = 6\]

OpenStudy (anonymous):

thank u...

OpenStudy (anonymous):

same answer and work - 16-10=6